# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v2

@@ VI Simulation results
Each data point in the plots is based on at least 100 logical error events in the low error rate regime. At LER $10^{-6}$, this represents an error bar less than $2\times 10^{-7}$, which is roughly the size of the markers for the data points. Thus the error bars are neglected for clarity.
We observe that both parallel and serial schedules in MBP4+ADOSD4 yield similar results for toric codes and surface codes. However, the parallel schedule outperforms the serial schedule on the [[882, 48, 16]] GHP code. We believe that this is because that the parallel schedule typically results in unsettled oscillations in BP, which can be resolved by OSD, whereas the serial schedule may make premature decisions on some bits, leading to incorrect convergence in the decoding process. Therefore, we consider the parallel schedule for the following simulations of MBP4+ADOSD4. Similar observations are also made in [55, 65].
decoder source note
MBP4+ADOSD4 based on Algorithm 4
MBP4+OSD4-$w$ this paper reliability metric based on Definition 4
MBP4+mOSD4-$w$ reliability metric based on Definition 5
AMBP4 [32] No postprocessing. Can act as
a predecoder for ADOSD4 and OSD4
BP+mOSD($w$,$\lambda$) [43] binary BP+ order-w OSD4
with additional flips on $\lambda$ positions
[…]
a greedy search on (${\lambda\choose 2}$ positions
BP+GDG [55] binary BP with guided decimation
guessing based on a decision tree
TABLE I: Various decoders compared in this paper.
For comparison, we consider several decoders listed in Table I. Except for AMBP, we do not implement these decoders ourselves but instead compare our results with those reported in the literature.
@@
In [66], it was shown that a random order schedule improves BP decoding performance. Here, we implement the AMBP decoder with a random order serial schedule, achieving better performance than the results reported in [32]. Note that AMBP incorporates various techniques for performance enhancement, and we do not focus on optimizing its performance in the following simulations.
VI-A Optimization of the parameter $\alpha$ in MBP4 for ADOSD.
@@
In this subsection, we demonstrate how to optimize the parameter $\alpha$ of our MBP4+ADOSD4 scheme using LCS codes [60] with ${\mathrm{T}}=100$ and $\theta=0.999995$. For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance. However, for degenerate codes, selecting $\alpha>1$ can result in more stable message passing, leading to improved OSD performance. This optimization should be conducted for each code and potentially for each error rate, though a single value of $\alpha$ often applies across an entire code family.
@@
LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions. Decoding LCS codes presents the most challenging problem for AMBP4 among all the quantum codes considered in this paper. We find that the AMBP4 decoder [32] is ineffective for LCS codes, despite its superior performance over binary BP-OSD decoders [43, 44] on all other codes.
@@
Figure 2: MBP4+ADOSD4 decoding with various values of $\alpha$ for the $[[175,7,7]]$ LCS code. The curve MLE is taken from [60]. The notation $\alpha\in(1.6,-0.01,0.5)$ means that $\alpha$ is tested in the sequence $1.60,1.59,1.58,\dots,0.51,0.50$.
@@
Figure 2 compares various decoders on the $[[175,7,7]]$ LCS code, including an MLE (independent X/Z) reference curve from [60]. This MLE curve serves as a benchmark, as a decoder capable of correcting errors of weight up to $3$ should exhibit a performance curve parallel to it at low error rates.
@@
In [60], the BP-OSD-CS60 decoder was shown to closely match the MLE curve on this code. Since the decoding of $X$ and $Z$ errors is performed independently, the MLE curve represents optimal performance under the assumption that $X$ and $Z$ errors are uncorrelated. As MBP4+ADOSD4 leverages the $X/Z$ correlations, it significantly outperforms both the MLE curve and BP-OSD-CS60 by more than 15 times. In contrast, AMBP4 encounters an error floor for $\epsilon\leq 0.03$.
@@
For improved performance, $\alpha$ can be adaptively chosen for each $\epsilon$, as demonstrated in Section VI-D.
VI-B Robustness of the maximum iterations in MBP4+ADOSD4
Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations ${\mathrm{T}}$ for the $[[175,7,7]]$ LCS code.
Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. We note that in ADOSD4, OSD-$w$ is invoked when HRSR fails. As a result, the overall algorithm maintains high accuracy. However, the failures of the HRSR algorithm can impact the simulation time, as OSD-$w$ is significantly more time-consuming.
@@ VI-B Robustness of the maximum iterations in MBP4+ADOSD4
TABLE II: The average number of iterations for MBP4 decoding on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=200$.
Table II presents the average number of iterations for ${\mathrm{T}}=200$. It can be observed that the average number of iterations rapidly decreases to one or two when $\epsilon<0.01$. Since we are targeting lower error rates, to balance accuracy and efficiency, we typically choose ${\mathrm{T}}=100$.
VI-C Selection of Reliability Threshold
We now discuss the selection of the reliability threshold. Figure 4 presents an example for the $[[175,7,7]]$ LCS code, which is representative of our simulation results.
Based on our experience, at high error rates, the reliability threshold has little impact on performance. However, at low error rates, where higher accuracy is required, the choice of $\theta$ depends on both ${\mathrm{T}}$ and the specific code. For instance, MBP4+ADOSD4 achieves excellent performance on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=2$ and $\theta=0.5$ but performs poorly on the (4.8.8) color code with $d=5$ under the same setting.
With ${\mathrm{T}}=100$ and $\theta=0.999995$, we generally observe good and stable decoding performance across all codes considered, as shown in Figure 4. Therefore, we adopt these values for our analysis in the following subsections.
Figure 4: MBP4+ADOSD4 decoding with various values of reliability threshold $\theta$ for the $[[175,7,7]]$ LCS code.
VI-D Generalized hypergraph-product codes
The $[[882,48,16]]$ GHP code [43] is highly degenerate, which causes conventional BP to perform poorly. Figure 5 presents the performance curves of different MBP and OSD combinations, where $\alpha$ is either fixed at 1.6 or adaptively chosen as $\alpha(\epsilon)$.
For comparison, we also include the performance of AMBP4 [32] and BP+mOSD($w$, $\lambda$[43]. BP+mOSD($w$, $\lambda$) exhibits an early error floor around $\epsilon=0.09$. Both MBP4+ADOSD4 and MBP4+OSD4-2 with $\alpha=1.6$ outperform BP+mOSD($15$, $15$[43] by more than an order of magnitude at $\epsilon=0.05$.
@@ VI-D Generalized hypergraph-product codes
Figure 5: Comparison of various decoders on the $[[882,48,16]]$ GHP code. The curve BP+mOSD(15,15) is from [43].
Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code with $\alpha=1.6$.
In ADOSD4, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4). Herein we also simulate the performances of MBP4+mOSD based on the reliability metrics of Entropy or Max as shown in Figure 6, including mOSD4  as well. Two conclusions are as follows:
  1. 1.
    The reliability metric in Definition 4, which incorporates historical message information, outperforms other metrics that rely solely on the last belief distributions.
  2. 2.
    The performances of mOSD based on Definition 5, Entropy, or Max are mostly comparable; mOSD-$w$ based on Definition 5 is slightly better than the other two metrics at lower error rate for $w=2$.
VI-E Bivariate bicycle codes
BB codes are a class of QLDPC codes proposed in [56], featuring a high error threshold and a significantly higher code rate than topological codes. Several attempts at decoding this code family have been presented in [54, 55] under the code capacity noise model. However, the BP+CB decoder in [54] does not perform as well as BP-OSD-0. The BP+GDG decoder in [55] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [55, Figure 4].
Figure 7 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than GDG decoding. For instance, consider the [[144,12,12]] BB code. At an $X$ error rate of $0.02$ (corresponding to a depolarizing rate of about $0.03$), BP+GDG achieves an LER of approximately $10^{-4}$. In contrast, both MBP4+ADOSD4 and AMBP4 achieve an LER below $4\times 10^{-6}$ at the same depolarizing rate. Moreover, MBP4+ADOSD4 outperforms AMBP4.
Figure 7: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes.
VI-F 2D topological codes
Figure 8: MBP4+OSD4-2, MBP4+ADOSD4, and AMBP4 decoding performance for various $[[d^{2},1,d]]$ rotated surface codes.
Figure 9: (a) The threshold of MBP4+OSD4-2 on the $[[d^{2},1,d]]$ rotated surface codes with $\alpha=1.0$ is about $17.67\%$ from the simulations. For each data point, we collected at least 10,000 logical error events. The dashed line stands for no error correction. (b) The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach. The estimation is based on data from $d=7$ to $d=17$, ensuring that the results are not influenced by biased information.
Figure 8 presents the performance curves of MBP4+ADOSD4 and MBP4+OSD4-2 on $[[d^{2},1,d]]$ rotated surface codes with distances ranging from 5 to 17. They exhibit excellent performance, closely following the dotted reference curves.
The threshold analysis for $[[d^{2},1,d]]$ rotated surface codes, conducted using the scaling ansatz method [67, 68], is shown in Figure 9. The results show a threshold of approximately 17.67% for BP4+OSD4-2, with MBP4+ADOSD4 exhibiting nearly identical performance. Notably, they outperform AMBP4, particularly for larger $d$, resulting in a higher threshold compared to the  16% threshold of AMBP4.
Similar analyses are performed for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes using both BP4+OSD4-2 and BP4+ADOSD4. The summarized threshold results are presented in Table III. ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes.
Figure 10: AMBP4 and MBP4+ADOSD4 decoding performance for various (4.8.8) color codes.
TABLE III: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
@@ VI-E Bivariate bicycle codes
The thresholds for the color codes are lower than those of toric and surface codes. This is evident from the performance curves of the (4.8.8) color codes as shown in Figure 10. While the slopes of these curves increase as the error rate decreases, they do not closely match the slopes of the dotted reference curves. This suggests that some low-weight errors, even those with weight less than $t+1$, may remain uncorrected by these decoders. Consequently, additional techniques or higher-order OSD may be necessary.
TABLE IV: Comparison of the time consumption for surface codes at LER about $10^{-6}$.
distance failure rate iteration ($\mu$s) ADOSD4 ($\mu$s) OSD4-2 ($\mu$s) when MBP4 succeeds in MBP4 iterations
@@
VI-G Time consumption
In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes. We fix $\alpha=1$ in MBP4.
VI-G1 Time Consumption of ADOSD4 vs. OSD4-2 and BP
Table IV lists the average time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15 at LER arouond $10^{-6}$. The unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer. To achieve LER $10^{-6}$ using MBP4+ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively. At least $10^{5}$ samples are collected for each case.

### v3

@@ VII Simulation results
Each data point in the plots is based on at least 100–1000 logical error events in the low error rate regime. At LER $10^{-6}$, this represents an error bar less than $2\times 10^{-7}$, which is roughly the size of the markers for the data points. Thus the error bars are neglected for clarity.
@@
We use only the parallel schedule for the following simulations of MBP4+ADOSD4. In contrast, [83] showed that a random-order schedule can improve BP decoding performance. Accordingly, we implement the AMBP decoder with a random-order serial schedule, achieving better performance than the results reported in [32].
In ADOSD4 simulations, $\Gamma$ is chosen according to (17).
VII-A Parameter Selection and Tuning
We use the MBP4 decoder introduced in [32] as a predecoder for OSD. MBP4 features a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced. For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [13]. We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used. The optimization of $\alpha$ is discussed in Section VII-A1.
The maximum number of BP iterations, $\mathrm{T}$}, is set to $100$ and the soft reliability threshold to $\theta=0.999995$ for simulations under the code-capacity noise model to ensure high accuracy, while for the circuit-level noise model, $\mathrm{T}$} is set to $10$ and $\theta=0.99$ to prioritize computational efficiency.
VII-A1 Optimization of the parameter $\alpha$ in MBP4 for ADOSD
@@
In this subsection, we demonstrate how to optimize the parameter $\alpha$ of our MBP4+ADOSD4 scheme using LCS codes [57]. For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance. However, for degenerate codes, selecting $\alpha>1$ can result in more stable message passing, leading to improved OSD performance. This optimization should be conducted for each code and potentially for each error rate, though a single value of $\alpha$ often applies across an entire code family.
@@
LCS codes are a specialized type of quantum codes that combines the features of surface codes and lift-product constructions. Decoding LCS codes presents the most challenging problem for AMBP4 among all the quantum codes considered in this paper. We find that the AMBP4 decoder [32] is ineffective for LCS codes, despite its superior performance over binary BP-OSD decoders [44, 51] on all other codes.
@@
Figure 2: MBP4+ADOSD4 decoding with various values of $\alpha$ for the $[[175,7,7]]$ LCS code. The curve MLE is taken and rescaled from [57]. The notation $\alpha\in(1.6,-0.01,0.5)$ means that $\alpha$ is tested in the sequence $1.60,1.59,1.58,\dots,0.51,0.50$ with ${\mathrm{T}}=100$ and $\theta=0.999995$.
@@
Figure 2 compares various decoders on the $[[175,7,7]]$ LCS code, including an MLE (independent X/Z) reference curve from [57]. To compare with the MLE curve under $X$-only noise, we rescale the physical error rate by a factor of $3/2$ so that the effective $X$ error probability matches that under depolarizing noise. This MLE curve serves as a benchmark, as a decoder capable of correcting errors of weight up to $3$ should exhibit a performance curve parallel to it at low error rates.
@@
In [57], the BP-OSD-CS60 decoder [51] was shown to closely match the MLE curve on this code. Since the decoding of $X$ and $Z$ errors is performed independently, the MLE curve represents optimal performance under the assumption that $X$ and $Z$ errors are uncorrelated. As MBP4+ADOSD4 leverages the $X/Z$ correlations, it significantly outperforms both the MLE curve and BP-OSD-CS60 by more than 15 times. In contrast, AMBP4 encounters an error floor for $\epsilon\leq 0.03$.
@@
VII-A2 Robustness of the maximum iterations in MBP4+ADOSD4
Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations $\mathrm{T}$} for the $[[175,7,7]]$ LCS code with $\theta=0.999995$.
Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. We note that in ADOSD4, OSD-$w$ is invoked when RSR fails. As a result, the overall algorithm maintains high accuracy.
@@ VII-A2 Robustness of the maximum iterations in MBP4+ADOSD4
TABLE I: The average number of iterations for MBP4 decoding on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=200$.
Table I presents the average number of iterations for ${\mathrm{T}}=200$. It can be observed that the average number of iterations rapidly decreases to one or two when $\epsilon<0.01$. Since we are targeting lower error rates, to balance accuracy and efficiency, we typically choose ${\mathrm{T}}=100$.
VII-B Code capacity noise model decoding simulations
VII-B1 Bivariate bicycle codes
BB codes are a class of QLDPC codes proposed in [56], featuring a high error threshold and a significantly higher code rate than topological codes. Several attempts at decoding this code family have been presented in [52, 53] under the code capacity noise model. However, the BP+CB decoder in [52] does not perform as well as BP-OSD-0. The BP+GDG decoder in [53] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [53, Figure 4].
Figure 4 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than BP+GDG decoding [53]. For instance, consider the [[144,12,12]] BB code. At an $X$ error rate of $0.02$ (corresponding to a depolarizing rate of about $0.03$), BP+GDG achieves an LER of approximately $10^{-4}$. In contrast, both MBP4+ADOSD4 and AMBP4 achieve an LER below $4\times 10^{-6}$ at the same depolarizing rate. Moreover, MBP4+ADOSD4 outperforms AMBP4.
Figure 4: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes with ${\mathrm{T}}=100$ and $\theta=0.999995$.
VII-B2 2D topological codes
Figure 5: MBP4+OSD4-2, MBP4+ADOSD4, and AMBP4 decoding performance for various $d^{2},1,d]]$ rotated surface codes with ${\mathrm{T}}=100$ and $\theta=0.999995$.
Figure 6: (a) The threshold of MBP4+OSD4-2 on the $d^{2},1,d]]$ rotated surface codes with $\alpha=1.0$, $T=100$, and $\theta=0.999995$ is about $17.67\%$ from the simulations. For each data point, we collected at least 10,000 logical error events. The dashed line stands for no error correction. (b) The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach. The estimation is based on data from $d=7$ to $d=17$, ensuring that the results are not influenced by biased information.
Figure 5 presents the performance curves of MBP4+ADOSD4 and MBP4+OSD4-2 on $d^{2},1,d]]$ rotated surface codes with distances ranging from 5 to 17. They exhibit excellent performance, closely following the dotted reference curves.
The threshold analysis for $d^{2},1,d]]$ rotated surface codes, conducted using the scaling ansatz method [84, 85], is shown in Figure 6. The results show a threshold of approximately 17.67% for BP4+OSD4-2, with MBP4+ADOSD4 exhibiting nearly identical performance. Notably, they outperform AMBP4, particularly for larger $d$, resulting in a higher threshold compared to the  16% threshold of AMBP4.
Similar analyses are performed for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes using both BP4+OSD4-2 and BP4+ADOSD4. The summarized threshold results are presented in Table II. ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes.
TABLE II: Thresholds for 2D topological codes under depolarizing errors. The parameters are $\alpha=1.0$, $T=100$, and $\theta=0.999995$
@@ VII-B1 Bivariate bicycle codes
TABLE III: Comparison of the time consumption for surface codes at LER about $10^{-6}$ with $\alpha=1.0$, $T=100$, and $\theta=0.999995$.
@@
VII-B3 Time Consumption of ADOSD4 vs. OSD4-2 and BP
[…]